OpenStudy (gelow):

What is 3 2/5* 2 1/3 ? Express the answer in simplest form.

OpenStudy (gelow):

@Catlover5925

OpenStudy (anonymous):

We need to change the mixed numbers into improper fractions. Do you know how to do that?

OpenStudy (anonymous):

@gelow

OpenStudy (gelow):

yeah it would be 17/5 * 7/3

OpenStudy (gelow):

@AnswerMyQuestions

OpenStudy (anonymous):

Yes, so just multiply straight across.

OpenStudy (anonymous):

\(\Huge\frac{17\ \times\ 7}{5\ \times\ 3}\)

OpenStudy (gelow):

ok thank you here is a medal

OpenStudy (gelow):

i got it

OpenStudy (anonymous):

No problem! :)